# Evaluating the Dynamics of Membership Privacy in Deep Learning

### v1

@@ Introduction
Figure 1: An illustration of sample-level vulnerability trajectories. Each line tracks a sample’s vulnerability path, $\tau(z)=(v_{1}(z),\dots,v_{*}(z))$, over * training epochs. Each state $v_{t}(z)$ is estimated from $N$ model variations at epoch $t$. The diverse paths reveal how membership encoding is influenced by sample properties and training dynamics—a process obscured by post-hoc analysis that captures only the final state ( ).
@@
    We reveal a strong and consistent correlation between a sample’s intrinsic learning difficulty, quantified by metrics like influence functions, learning iterations, and epistemic uncertainty, and its ultimate susceptibility to MIAs. Crucially, we demonstrate that the vulnerability of these high-risk, “hard-to-learn” samples is exposed within a critical early-training window, paving the way for proactive and targeted defenses.
@@ A Formalization for Per-Sample Vulnerability
Table 2: Metrics on different models (subscripts indicate standard deviation; values with scaled by $10^{-3}$).. Values ×$\times 10^{-3}$.
Metric cnn32-3-max cnn64-3-max wrn28-2
Motion
Speed ($\bar{v}$) 1.100.68 2.851.20 2.951.30
@@
Table 3: Correlation between sample hardness metrics and membership vulnerability. Pearson correlation coefficients are shown for both static vulnerability and dynamic vulnerability. Higher absolute values indicate stronger predictive power.
Hardness Metric Target CNN-32 CNN-64 WRN-28
@@
Low-level Optimization Signal
Gradient Norm $\alpha$ 0.000 0.005 -0.012 0.003 0.004 -0.012 0.002 0.001 0.001
$v_{\alpha}$ 0.000 0.007 -0.005 0.003 0.005 -0.008 0.004 0.008 0.002
Cumulative Learning Effort
Iteration Learned $\alpha$ 0.411 0.708 -0.613 0.559 0.755 -0.624 0.564 0.692 -0.493
$v_{\alpha}$ 0.402 0.674 -0.593 0.548 0.734 -0.615 0.565 0.719 -0.554
Influence Function $\alpha$ 0.450 0.680 -0.600 0.510 0.740 -0.630 0.480 0.710 -0.460
$v_{\alpha}$ 0.425 0.690 -0.580 0.500 0.720 -0.640 0.520 0.740 -0.520
Model Uncertainty
Aleatoric (data) $\alpha$ 0.500 0.778 -0.659 0.654 0.812 -0.647 0.779 0.885 -0.642
$v_{\alpha}$ 0.488 0.742 -0.619 0.635 0.786 -0.633 0.776 0.913 -0.764
Epistemic (model) $\alpha$ 0.505 0.766 -0.633 0.656 0.801 -0.612 0.808 0.908 -0.609
$v_{\alpha}$ 0.494 0.737 -0.591 0.636 0.777 -0.597 0.790 0.916 -0.725
Description. To understand why certain individual samples are more vulnerable than others, we investigate the link between the intrinsic characteristics of a data sample and its susceptibility to membership inference, demonstrating that a sample’s “hardness”—the difficulty a model has in learning it—is a strong predictor of its privacy risk. We quantify sample hardness using a set of metrics grouped into three categories: low-level optimization signals, cumulative learning effort, and model uncertainty. We then measure the Pearson correlation between these hardness metrics and two facets of privacy risk derived from our framework: 1) the final, static membership advantage ($\alpha$) after training, and 2) the dynamic membership encoding speed ($v_{\alpha}$), which captures the rate of change in advantage during training. The results, compiled in Table 3, are presented across multiple architectures and are conditioned on a sample’s final vulnerability status (i.e., sample-level membership advantage $\alpha>0$ or $\alpha\leq 0$).
Results. Table 3 summarizes the correlation results, highlighting a clear hierarchy in the predictive strength of different hardness metrics. Notably, there is a stark difference between metrics capturing long-term learning behavior and those representing instantaneous updates. Low-level optimization signals, such as gradient norm, show negligible correlation with both static vulnerability (membership advantage $\alpha$ after training) and dynamic vulnerability (encoding speed $v_{\alpha}$ during training). This result underscores that privacy leakage is not driven by isolated, large parameter updates but rather emerges from the cumulative learning trajectory of individual samples.
Conversely, metrics representing the cumulative algorithmic effort required to learn a sample—such as the iteration learned and influence function—exhibit strong positive correlations with privacy risk. This relationship is even more pronounced for model uncertainty metrics. Epistemic uncertainty, which quantifies the model’s inherent uncertainty about a sample, emerges as an exceptionally strong predictor. For the WRN-28 model, the correlation between epistemic uncertainty and dynamic vulnerability ($v_{\alpha}$) for high-risk samples reaches a notably high value of 0.916. This indicates that the model struggles to generalize these difficult samples, resorting instead to memorization, and thus leaving a substantial information-theoretic footprint.
@@
The analysis becomes even more insightful when conditioning on the final vulnerability status. For the vulnerable subset ($\alpha>0$), the positive correlations are consistently strong, indicating that among samples the model fails to protect, the extent of privacy leakage directly increases with learning difficulty. Conversely, within the non-vulnerable subset ($\alpha\leq 0$), the observed negative correlation suggests that samples characterized by high uncertainty or influence tend to be resolved through generalization rather than memorization, thus reducing privacy risk. Although this reveals a nuanced behavior within the subsets, the overall correlation across all samples remains positive, driven primarily by the strong positive relationship between sample hardness and vulnerability in the vulnerable subgroup.
(a) MNIST
(b) F-MNIST
(c) CIFAR10
(d) CINIC10
Figure 5: Sample-level membership advantage variation based on epochs using different datasets. The z-axis and color both represent the metric score of aleatoric uncertainty.
@@ Appendix B Consistency of Findings with Alternative MIA Methods
Table 4: Performance metrics across datasets using the  Shokri et al. (2017) attack for comparison with Table 1. The trends are consistent with LiRA, but absolute values are lower, reflecting a weaker attack. Values ×$\times 10^{-3}$.
Metric MNIST F-MNIST CIFAR10 CINIC10
Motion
Speed ($\bar{v}$) 0.250.30 0.950.61 2.151.10 2.801.15
@@
Information
Entropy ($H$) 4.380.06 4.250.40 4.610.72 4.750.55
$\Delta H$0.05 1.10 2.95 3.68
Dir. Angle -0.65 0.98 0.88 0.92
Clustering
Avg. ($\bar{k}$) 325030 2410350 1890400 1650380
@@
Table 5: Metrics on different models using the  Shokri et al. (2017) attack for comparison with Table 2. The trends are consistent with LiRA. Values ×$\times 10^{-3}$.
Metric cnn32-3-max cnn64-3-max wrn28-2
Motion
Speed ($\bar{v}$) 0.800.55 2.051.00 2.151.10
@@
Dir. Angle ($\theta$) -0.42 0.85 0.88
Clustering
Avg. ($\bar{k}$) 2250150.5 1815380.1 1890400
@@
First, the post-training distributions on the Vulnerability Plane exhibit the same structural changes as we seen in Figure 2. As shown in Figure 89, and 10, while the  Shokri et al. (2017) attack yields generally lower TPR and higher FPR values (shifting the entire distribution down and to the right), the core trend is preserved. For simpler datasets like MNIST, the sample population remains tightly clustered along the non-vulnerable diagonal. For complex datasets like CINIC-10, the distribution disperses significantly, populating the upper regions of the plane and confirming a more heterogeneous risk landscape.

### v2

@@ Introduction
Figure 1: An illustration of sample-level vulnerability trajectories. Each dot on a path represents a sample’s vulnerability state, characterized by its membership attack metric (FPR, TPR), estimated from $N$ model variations at a given epoch. The distinct paths illustrate how membership encoding evolves uniquely for each sample—nuances obscured by traditional post-hoc analyses that only assess the final state ( ).
@@
    We reveal a strong correlation between intrinsic sample hardness—quantified through metrics such as influence functions, learning iterations, and epistemic uncertainty—and membership vulnerability. Crucially, our analysis demonstrates that vulnerability of these “hard-to-learn” samples is predominantly determined early during training, indicating a critical window for proactive intervention. Moreover, we show that optimization methods explicitly promoting generalization, notably Sharpness-Aware Minimization (SAM) (Foret et al. 2021), substantially mitigate the encoding of membership information throughout training.
@@ A Formalization for Per-Sample Vulnerability
Table 2: Metrics on different models (subscripts indicate standard deviation; values with scaled by $10^{-3}$).
Group Metric cnn32-3-max cnn64-3-max wrn28-2
Motion Speed ($\bar{v}$) 1.100.68 2.851.20 2.951.30
@@
Table 3: Correlation between sample hardness metrics and membership vulnerability. Pearson correlation coefficients are shown for both static vulnerability ($\alpha$) and dynamic vulnerability ($v_{\alpha}$). Bold values denote $|r|\geq 0.80$ (very strong correlation).
Group Metric Target CNN-32 CNN-64 WRN-28
@@
Low-level Opt. Gradient Norm $\alpha$ 0.000 0.005 0.012-0.012- 0.012 0.003 0.004 0.012-0.012- 0.012 0.002 0.001 0.001
$v_{\alpha}$ 0.000 0.007 0.005-0.005- 0.005 0.003 0.005 0.008-0.008- 0.008 0.004 0.008 0.002
Cumulative Effort Iteration Learned $\alpha$ 0.411 0.708 0.613-0.613- 0.613 0.559 0.755 0.624-0.624- 0.624 0.564 0.692 0.493-0.493- 0.493
$v_{\alpha}$ 0.402 0.674 0.593-0.593- 0.593 0.548 0.734 0.615-0.615- 0.615 0.565 0.719 0.554-0.554- 0.554
Influence Function $\alpha$ 0.450 0.680 0.600-0.600- 0.600 0.510 0.740 0.630-0.630- 0.630 0.480 0.710 0.460-0.460- 0.460
$v_{\alpha}$ 0.425 0.690 0.580-0.580- 0.580 0.500 0.720 0.640-0.640- 0.640 0.520 0.740 0.520-0.520- 0.520
Uncertainty Aleatoric $\alpha$ 0.500 0.778 0.659-0.659- 0.659 0.654 0.812 0.647-0.647- 0.647 0.779 0.885 0.642-0.642- 0.642
$v_{\alpha}$ 0.488 0.742 0.619-0.619- 0.619 0.635 0.786 0.633-0.633- 0.633 0.776 0.913 0.764-0.764- 0.764
Epistemic $\alpha$ 0.505 0.766 0.633-0.633- 0.633 0.656 0.801 0.612-0.612- 0.612 0.808 0.908 0.609-0.609- 0.609
$v_{\alpha}$ 0.494 0.737 0.591-0.591- 0.591 0.636 0.777 0.597-0.597- 0.597 0.790 0.916 0.725-0.725- 0.725
Description. To understand why certain individual samples are more vulnerable than others, we examine the relationship between a sample’s intrinsic “hardness”—the difficulty a model encounters when learning it—and susceptibility to membership inference. We quantify hardness using three categories of metrics. First, low-level optimization signals, such as Gradient Norm (Chen et al. 2018), indicate the magnitude of parameter updates per sample. Second, cumulative learning effort metrics include Iteration Learned (Toneva et al. 2019), which tracks how long a sample takes to be consistently learned, and Influence Functions (Koh and Liang 2020), measuring a sample’s impact on model predictions. Third, model uncertainty metrics, specifically Aleatoric and Epistemic Uncertainty (Hüllermeier and Waegeman 2021), assess the model’s confidence or lack thereof, with high epistemic uncertainty typically linked to samples challenging to generalize and thus at greater risk of memorization.
We compute the Pearson correlation between these hardness metrics and two dimensions of privacy risk: the static membership advantage ($\alpha$) after training, and the dynamic membership encoding speed ($v_{\alpha}$) that quantifies how rapidly privacy vulnerability evolves.
Results. Table 3 summarizes correlation findings, revealing distinct predictive power across different hardness metrics. Low-level optimization signals like gradient norm exhibit negligible correlations with both static and dynamic vulnerability, indicating privacy leakage arises from cumulative learning trajectories rather than isolated parameter updates.
In contrast, cumulative learning effort metrics—iteration learned and influence functions—show strong positive correlations with privacy risk. Model uncertainty metrics, especially epistemic uncertainty, demonstrate an exceptionally high correlation with vulnerability (0.916 for dynamic vulnerability in the WRN-28 model). This indicates that high epistemic uncertainty, reflecting the model’s struggle to generalize, significantly increases memorization risk.
@@
Conditioning on the final vulnerability status offers additional insights. For the vulnerable subset ($\alpha>0$), the positive correlation between learning difficulty and privacy risk is particularly pronounced, emphasizing that harder-to-learn samples incur greater leakage. Conversely, within the non-vulnerable subset ($\alpha\leq 0$), negative correlations suggest that challenging samples in this group enhance model generalization rather than memorization, reducing privacy risks. Despite this nuanced subset behavior, the overall correlation remains positive, primarily driven by the strong relationship observed in the vulnerable subgroup.
@@ Appendix B Consistency of Findings with Alternative MIA Methods
Table 4: Performance metrics across datasets using the Shokri et al. (2017) attack (subscripts indicate standard deviation; values with scaled by $10^{-3}$).
Group Metric MNIST F-MNIST CIFAR10 CINIC10
Motion Speed ($\bar{v}$) 0.250.30 0.950.61 2.151.10 2.801.15
@@
Information Entropy ($H$) 4.380.06 4.250.40 4.610.72 4.750.55
$\Delta H$ 0.05-0.05- 0.05 1.10 2.95 3.68
Dir. Angle ($\theta$) 0.65-0.65- 0.65 0.98 0.88 0.92
Clustering Avg. ($\bar{k}$) 325030 2410350 1890400 1650380
@@
Table 5: Metrics on different models using the Shokri et al. (2017) attack (subscripts indicate standard deviation; values with scaled by $10^{-3}$).
Group Metric cnn32-3-max cnn64-3-max wrn28-2
Motion Speed ($\bar{v}$) 0.800.55 2.051.00 2.151.10
@@
Dir. Angle ($\theta$) 0.42-0.42- 0.42 0.85 0.88
Clustering Avg. ($\bar{k}$) 2250150.5 1815380.1 1890400
@@
First, the post-training distributions on the Vulnerability Plane exhibit the same structural changes as we seen in Figure 2. As shown in Figure 78, and 9, while the  Shokri et al. (2017) attack yields generally lower TPR and higher FPR values (shifting the entire distribution down and to the right), the core trend is preserved. For simpler datasets like MNIST, the sample population remains tightly clustered along the non-vulnerable diagonal. For complex datasets like CINIC-10, the distribution disperses significantly, populating the upper regions of the plane and confirming a more heterogeneous risk landscape.